OpenStudy (anonymous):

find the derivative of f(x)=(x-3)(x+2)^3

OpenStudy (anonymous):

Using formula : \[(uv)'= u'v+v'u\]. Do it carefully.

OpenStudy (anonymous):

do i put the exponent in front of the (x-3) or (x+2)?

OpenStudy (anonymous):

for this function, you should think that : \[u(x)=x-3\] and \[v(x)=(x+2)^3\] Then using formula above

OpenStudy (anonymous):

oh okay so i got \[1\times(x+2)^3+3(1)^2(x-3)\] is this correct?

OpenStudy (anonymous):

No, you should study next formula : \[(h^n (x))'=n.h^{n-1}.h'\] Do v'(x) again.

OpenStudy (anonymous):

\[(h^n(x))'=n.h^{n−1} (x).h'(x)\]

OpenStudy (anonymous):

so v'(x)=3(x+2)^2???

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so therefore would the next step be (x+2)^3+3(x+2)^2(x-3)?

OpenStudy (anonymous):

Yup, that's the derivative. :-)

OpenStudy (anonymous):

You can simplify it if you want

OpenStudy (anonymous):

wouldn't that be x^3+15x^2+25x+9?

OpenStudy (anonymous):

\[(x+2)^2(4x-7)\]

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

\[(x+2)^3+3(x+2)^2(x-3)=(x+2)^2 (x+2+3(x-3))=...\] it's here.